"solve:4w-3w+3=-4+7"
Answer by Flannery(124)

First, 4w-3w=w, so
w+3=-4+7
then -4+7=7-4=3, so
w+3=3
then subtract 3 from both sides to isolate w, so
w=0
4(0)-3(0)+3=-4+7
3=-4+7
3=3